OpenStudy (joftheworld):

I need help plotting three points "whole numbers" on a linear equation.. -3y=2x-7

OpenStudy (anonymous):

just choose three arbitrary points for x: say x=-1,x=0 and x=1; x=-1=> -3y=2*(-1)-7 so y=3 (-1,3); x=0=> -3y=2*0-7 so y= 7/3 (0,7/3) and x=1 => -3y=2*1-7=5/3 (1, 5/3) now mark all these points on the coordinate system and connect the dots

OpenStudy (joftheworld):

I need whole numbers..

OpenStudy (anonymous):

ok then we should know how to arrange them.. now: when you plug in 0, you get whole numbers (-1,3); let's try x=-4, then -3y=2*(-4)-7 => y=5 now we have (-4,5); for the finl one, let's try x=2, then -3y=2*(2)-7 => y=1 so (2,1)

OpenStudy (joftheworld):

that worked thanks! do you mind helping me with 5 more ?